# On the geometry of spaces of filtrations on local rings

### v1

@@ 2. Preliminaries
In this section, we recall definitions and basic properties that will be needed in the paper. Several results in Section 2.2, including a characterization of saturated filtrations Propositon 2.20, the lattice structure Definition-Lemma 2.22 and the distributivity Proposition 2.24, seem to be new.
@@ 2.1.1. Filtrations
    (left-continuouity) $\mathfrak{a}_{\lambda}=\mathfrak{a}_{\lambda-\epsilon}$ for any $\lambda\in\mathbb{R}_{>0}$ and $0<\epsilon\ll 1$, and
@@
    (multiplicativity) $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ for any $\lambda,\mu\in\mathbb{R}_{>0}$.
@@
Let $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ be two $\mathfrak{m}$-filtrations. Their intersection is $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}:=(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})_{\lambda\in\mathbb{R}_{>0}}$. It is not hard to check that this is again an $\mathfrak{m}$-filtrations.
@@
for some $l\in\mathbb{Z}_{>1/\lambda}$, where the first equality follows from the left-continuouity of $\mathfrak{a}_{\bullet,i}$, and the last equality uses the fact that $R/\mathfrak{b}_{\lambda}$ is Artinian. This proves that condition (2) holds for $\{\cap_{i\in I}\mathfrak{a}_{\lambda,i}\}$, and thus it is an $\mathfrak{m}$-filtration.
@@
Given an $\mathfrak{m}$-primary ideal $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$, we have an $\mathfrak{m}$-filtration $\mathfrak{a}^{\bullet}$, the $\mathfrak{a}$-adic filtration, defined by $\{\mathfrak{a}^{\lceil\lambda\rceil}\}_{\lambda\in\mathbb{R}_{>0}}$. Thus we get an injection from the set $\mathcal{I}_{\mathfrak{m}}$ to the set of all $\mathfrak{m}$-filtrations by $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$ In particular, there is a canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}$.
@@
Unlike the intersection, it can be hard to determine the join of two general filtrations. In Section 2.2, we will define a similar notion, the saturated join, and give a formula for it. The following proposition, which is a combination of Definition-Lemma 2.2 and Defnition-Lemma 2.3, justifies the term join.
@@
The set of all $\mathfrak{m}$-filtrations with the partial order by inclusion form a lattice, where the meet of two elements is given by their intersection, and the join is defined as in Definition-Lemma 2.3. ∎
@@ 2.1.2. Norms and valuations
As in the global case, we adopt the additive terminology here. For a norm $\chi$ on $R$, $\|\cdot\|_{\chi}:=e^{-\chi(\cdot)}$ defines a non-Archimedean norm on $R$ in the sense of [BGR84].
@@
$\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f):=\sup\{\lambda\in\mathbb{R}_{>0}\mid f\in\mathfrak{a}_{\lambda}\},$ (2.1)
@@
$\mathfrak{a}_{\lambda}(\chi):=\{f\in R\mid\chi(f)\geq\lambda\}.$ (2.2)
@@
Conversely, let $\chi$ be an $\mathfrak{m}$-seminorm. We first show that for any $\lambda>0$, $\mathfrak{a}_{\lambda}(\chi)$ defined in (2.2) is an $\mathfrak{m}$-primary ideal. It is an ideal since $\chi(f-g)\geq\min\{\chi(f),\chi(-g)\}=\min\{\chi(f),\chi(g)\}$ and $\chi(fg)\geq\chi(f)+\chi(g)\geq\min\{\chi(f),\chi(g)\}$. Choose a set of generators $f_{1},\ldots,f_{r}$ of $\mathfrak{m}$, then for any $f\in\mathfrak{m}$ we have $\chi(f)\geq\chi(\mathfrak{m})\coloneqq\min_{1\leq i\leq r}\{\chi(f_{i})\}>0$. Let $C\coloneqq\lceil\lambda/\chi(\mathfrak{m})\rceil$, then $\chi(f^{C})\geq C\chi(f)\geq\lambda$, that is, $\mathfrak{m}^{C}\subset\mathfrak{a}_{\lambda}(\chi)$. Note that the above argument also shows that $\mathfrak{a}_{\lambda}(\chi)=\mathfrak{m}$ for $0<\lambda<\chi(\mathfrak{m})$. Clearly the collection of $\mathfrak{m}$-primary ideals $\{\mathfrak{a}_{\lambda}(\chi)\}$ is decreasing, and it is multiplicative by condition (3) of Definition 2.7. To show left continuity, note that we can write $\mathfrak{a}_{\lambda}(\chi)$ as an decreasing intersection
@@ 2.1.5. Multiplicity of a filtration
$\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}=\inf_{m}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}};$ (2.3)
@@ 2.1.6. Linearly boundedness
By Krull intersection theorem, a linearly bounded $\mathfrak{m}$-filtration satisfies $\cap_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}=\{0\}$. In particular, by Definitin-Lemma 2.8, we have the following
@@
An $\mathfrak{m}$-filtration is linearly bounded if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is a linearly bounded $\mathfrak{m}$-norm.
We will consider filtrations with positive multiplicity, which are exactly those which are linearly bounded by the following lemma.
@@
There is a partial order $\leq$ on $\mathcal{N}$ defined pointwisely on $R$, that is,
$\chi\leq\chi^{\prime}\text{ if and only if }\chi(f)\leq\chi^{\prime}(f)\text{ for any }f\in R,$
@@ 2.2. Saturated filtrations
Conversely, assume that $\mathfrak{a}_{\bullet}\nsubseteq\mathfrak{b}_{\bullet}$. Then there exists $\lambda\in\mathbb{R}_{>0}$ and $f\in\mathfrak{a}_{\lambda}$ such that $f\notin\mathfrak{b}_{\lambda}$. This implies that there exists $\epsilon>0$ and $v\in\mathrm{DivVal}_{X,x}$ such that $v(f)<\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon$. Hence for any $m\in\mathbb{Z}_{>0}$, we have
@@
In other words, $\widetilde{\mathfrak{a}}_{\bullet}$ is the unique maximal $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}\supset\mathfrak{a}_{\bullet}$ such that $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$. We now prove a formula for saturations.
@@
$\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}).$ (2.4)
@@
$\mathfrak{a}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a}_{\bullet})})$ (2.5)
@@
applying the formula (2.5) to $\widetilde{\mathfrak{a}}_{\bullet}$ and $\widetilde{\mathfrak{b}}_{\bullet}$ respectively, we know that $\mathfrak{a}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}\subset\mathfrak{c}_{\bullet}$ and similarly $\mathfrak{b}_{\bullet}\subset\mathfrak{c}_{\bullet}$. Hence $v(\mathfrak{c}_{\bullet})\leq\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$ for any $v\in\mathrm{Val}_{R,\mathfrak{m}}$. By (2.4), for any $v\in\mathrm{DivVal}_{X,x}$ we have $v(\mathfrak{c}_{\bullet})\geq v(\mathfrak{a}_{\bullet}(v/\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}))=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$. Thus
$v(\mathfrak{c}_{\bullet})=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}.$ (2.6)
We now verify that $\mathfrak{c}_{\bullet}\in\mathrm{Fil}$. Take $C\in\mathbb{R}_{>0}$ such that $\mathfrak{a}_{C\bullet}\subset\mathfrak{m}^{\bullet}$ and $\mathfrak{b}_{C\bullet}\subset\mathfrak{m}^{\bullet}$. Since $\mathfrak{m}^{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$, for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$, we have $Cv(\mathfrak{a}_{\bullet})\geq v(\mathfrak{m}^{\bullet})=v(\widetilde{\mathfrak{m}}^{\bullet})$ and similarly $Cv(\mathfrak{b}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$. By (2.6), $Cv(\mathfrak{c}_{\bullet})\geq v(\widetilde{\mathfrak{m}}^{\bullet})$. So by Lemma 2.16 $\mathfrak{c}_{\bullet}\subset\widetilde{\mathfrak{m}}^{\bullet}$ is linearly bounded.
@@
$\geq$ $\max\{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\},\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\}$
@@
where the second inequality follows from Proposition 2.23, and the last inequality is easy to check. By Proposition 2.23 again, we get $f\in(\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet}))_{\lambda}$. This shows that $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{c}_{\bullet}\cap\mathfrak{b}_{\bullet})\supset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})\cap\mathfrak{b}_{\bullet}$ and the equality must hold. ∎
@@
$(\widetilde{\mathfrak{a}^{\bullet}})_{\lambda}=\{f\in\mathfrak{m}\mid v(f)\geq\lambda\cdot v(\mathfrak{a}^{\bullet})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
@@
$(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\{f\in\mathfrak{m}\mid v(f)\geq mv(\mathfrak{a})=v(\mathfrak{a}^{m})\text{ for any }v\in\mathrm{DivVal}_{R,\mathfrak{m}}\}.$
@@
$\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$ (2.7)
@@ 2.3. Homogeneous filtrations
(2) Choose a set $\{f_{1},\ldots,f_{r}\}$ of generators of $\mathfrak{m}$ and $D\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{D}\subset\mathfrak{a}_{1}$. Then $f_{i}^{D}\in\mathfrak{a}_{1}$ and hence $f_{i}\in\mathfrak{a}_{1/D}$ by assumption. So $\mathfrak{m}\subset\mathfrak{a}_{1/D}$ and hence $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda<1/D$. The last assertion is straightforward by Definition-Lemma 2.8. ∎
@@
[CP24, Definition 3.5](c.f. [BFJ14, Section 4]) Given a filtration $\mathfrak{a}_{\bullet}$ on $R$, its asymptotic Samuel function $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is defined by $\widehat{\mathrm{ord}}_{\mathfrak{a}_{\bullet}}(f):=\lim_{k\to\infty}\frac{\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f^{k})}{k}$ for any $f\in R$.
Denote the corresponding filtration by $\widehat{\mathfrak{a}}_{\bullet}$, called the homogenization of $\mathfrak{a}_{\bullet}$.
The limit exists by Feteke’s lemma, see [CP24, Theorem 3.4]. Similar to Theorem 2.18 for the saturation, the homogenization has the following property.
@@
[CP24, Theorem 5.5] Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Then $\widehat{\mathfrak{a}}_{\bullet}$ is the unique largest $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\widehat{\mathrm{ord}_{\mathfrak{a}_{\bullet}}}=\widehat{\mathrm{ord}_{\mathfrak{b}_{\bullet}}}$.
[…]
@@
$\chi=\inf_{v\in I}v,$
where $I$ is a non-empty subset of $X^{\mathrm{an}}_{x}$.
@@
If $\chi(f)=\inf_{v\in I}v(f)$, then $\chi$ is homogeneous, since the same is true for each $v$.
The reverse implication is essentially a restatement of the Berkovich maximum modulus principle, applied to the seminormed ring $(R,\chi)$. See, for example, [Jon20, Theorem 2.2.1]

### v2

@@ 2. Preliminaries
In this section, we recall definitions and basic properties that will be needed in the paper. Several results in Section 2.2, including a characterization of saturated filtrations Proposition 2.20, the lattice structure Definition-Lemma 2.22 and the distributivity Proposition 2.24, seem to be new.
@@ 2.1.1. Filtrations
    (left-continuous) $\mathfrak{a}_{\lambda}=\mathfrak{a}_{\lambda-\epsilon}$ for any $\lambda\in\mathbb{R}_{>0}$ and $0<\epsilon\ll 1$, and
@@
    (multiplicative) $\mathfrak{a}_{\lambda}\cdot\mathfrak{a}_{\mu}\subset\mathfrak{a}_{\lambda+\mu}$ for any $\lambda,\mu\in\mathbb{R}_{>0}$.
@@
Let $\mathfrak{a}_{\bullet}$ and $\mathfrak{b}_{\bullet}$ be two $\mathfrak{m}$-filtrations. Their intersection is $\mathfrak{a}_{\bullet}\cap\mathfrak{b}_{\bullet}:=(\mathfrak{a}_{\lambda}\cap\mathfrak{b}_{\lambda})_{\lambda\in\mathbb{R}_{>0}}$. It is not hard to check that this is again an $\mathfrak{m}$-filtration.
@@
for some $l\in\mathbb{Z}_{>1/\lambda}$, where the first equality follows from the left-continuity of $\mathfrak{a}_{\bullet,i}$, and the last equality uses the fact that $R/\mathfrak{b}_{\lambda}$ is Artinian. This proves that condition (2) holds for $\{\cap_{i\in I}\mathfrak{a}_{\lambda,i}\}$, and thus it is an $\mathfrak{m}$-filtration.
@@
Given an $\mathfrak{m}$-primary ideal $\mathfrak{a}\in\mathcal{I}_{\mathfrak{m}}$, we have an $\mathfrak{m}$-filtration $\mathfrak{a}^{\bullet}$, the $\mathfrak{a}$-adic filtration, defined by $\{\mathfrak{a}^{\lceil\lambda\rceil}\}_{\lambda\in\mathbb{R}_{>0}}$. Thus we get an injection from the set $\mathcal{I}_{\mathfrak{m}}$ to the set of all $\mathfrak{m}$-filtrations defined by $\mathfrak{a}\mapsto\mathfrak{a}^{\bullet}$. In particular, there is a canonical $\mathfrak{m}$-filtration $\mathfrak{m}^{\bullet}$.
@@
Unlike the intersection, it can be hard to determine the join of two general filtrations. In Section 2.2, we will define a similar notion, the saturated join, and give a formula for it. The following proposition, which is a combination of Definition-Lemma 2.2 and Definition-Lemma 2.3, justifies the term join.
@@
The set of all $\mathfrak{m}$-filtrations with the partial order by inclusion forms a lattice, where the meet of two elements is given by their intersection, and the join is defined as in Definition-Lemma 2.3. ∎
@@ 2.1.2. Norms and valuations
Here, as in the global case, we adopt the additive terminology here. For a norm $\chi$ on $R$, $\|\cdot\|_{\chi}:=e^{-\chi(\cdot)}$ defines a non-Archimedean norm on $R$ in the sense of [BGR84].
@@
(2.1) $\mathrm{ord}_{\mathfrak{a}_{\bullet}}(f):=\sup\{\lambda\in\mathbb{R}_{>0}\mid f\in\mathfrak{a}_{\lambda}\},$
@@
(2.2) $\mathfrak{a}_{\lambda}(\chi):=\{f\in R\mid\chi(f)\geq\lambda\}.$
@@
Conversely, let $\chi$ be an $\mathfrak{m}$-seminorm. We first show that for any $\lambda>0$, $\mathfrak{a}_{\lambda}(\chi)$ defined in (2.2) is an $\mathfrak{m}$-primary ideal. It is an ideal since $\chi(f-g)\geq\min\{\chi(f),\chi(-g)\}=\min\{\chi(f),\chi(g)\}$ and $\chi(fg)\geq\chi(f)+\chi(g)\geq\min\{\chi(f),\chi(g)\}$. Choose a set of generators $f_{1},\ldots,f_{r}$ of $\mathfrak{m}$, then for any $f\in\mathfrak{m}$ we have $\chi(f)\geq\chi(\mathfrak{m})\coloneqq\min_{1\leq i\leq r}\{\chi(f_{i})\}>0$. Let $C\coloneqq\lceil\lambda/\chi(\mathfrak{m})\rceil$, then $\chi(f^{C})\geq C\chi(f)\geq\lambda$, that is, $\mathfrak{m}^{C}\subset\mathfrak{a}_{\lambda}(\chi)$. Note that the above argument also shows that $\mathfrak{a}_{\lambda}(\chi)=\mathfrak{m}$ for $0<\lambda<\chi(\mathfrak{m})$. Clearly the collection of $\mathfrak{m}$-primary ideals $\{\mathfrak{a}_{\lambda}(\chi)\}$ is decreasing, and it is multiplicative by condition (3) of Definition 2.7. To show left continuity, note that we can write $\mathfrak{a}_{\lambda}(\chi)$ as a decreasing intersection
@@ 2.1.3. Divisorial valuations
Recall that a valuation $v\in\mathrm{Val}_{R,\mathfrak{m}}$ defines a local ring $\mathcal{O}_{v}$, the valuation ring of $v$, by
$\mathcal{O}_{v}\coloneqq\{f\in\mathrm{Frac}(R)\mid v(f)\geq 0\}\text{ with the maximal ideal }\mathfrak{m}_{v}=\{f\in\mathcal{O}_{v}\mid v(f)\geq 0\}.$
The residue field of $v$ is $\kappa_{v}\coloneqq\mathcal{O}_{v}/\mathfrak{m}_{v}$.
@@ 2.1.5. Multiplicity of a filtration
(2.3) $\mathrm{e}(\mathfrak{a}_{\bullet})=\lim_{m\to\infty}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}}=\inf_{m}\frac{\mathrm{e}(\mathfrak{a}_{m})}{m^{n}};$
@@ 2.1.6. Linearly boundedness
By Krull intersection theorem, a linearly bounded $\mathfrak{m}$-filtration satisfies $\cap_{\lambda\in\mathbb{R}_{\geq 0}}\mathfrak{a}_{\lambda}=\{0\}$. In particular, by Definition-Lemma 2.8, we have the following
@@
An $\mathfrak{m}$-filtration is linearly bounded if and only if $\mathrm{ord}_{\mathfrak{a}_{\bullet}}$ is a linearly bounded $\mathfrak{m}$-norm.∎
We will consider filtrations with positive multiplicity, which are exactly the linearly bounded ones by the following lemma.
@@
There is a natural pointwise partial order $\leq$ on $\mathcal{N}$ on $R$, defined by
[…]
@@ 2.2. Saturated filtrations
Conversely, assume that $\mathfrak{a}_{\bullet}\nsubseteq\mathfrak{b}_{\bullet}$. Then there exist $\lambda\in\mathbb{R}_{>0}$ and $f\in\mathfrak{a}_{\lambda}$ such that $f\notin\mathfrak{b}_{\lambda}$. This implies that there exist $\epsilon>0$ and $v\in\mathrm{DivVal}_{X,x}$ such that $v(f)<\lambda\cdot v(\mathfrak{b}_{\bullet})-\epsilon$. Hence for any $m\in\mathbb{Z}_{>0}$, we have
@@
In other words, $\widetilde{\mathfrak{a}}_{\bullet}$ is the unique maximal $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}\supset\mathfrak{a}_{\bullet}$ such that $\mathrm{e}(\mathfrak{a}_{\bullet})=\mathrm{e}(\mathfrak{b}_{\bullet})$. We now prove a formula for saturations, which can be viewed as an analogue of the Berkovich maximal modulus principle, where one considers the subset of divisorial valuations of the Berkovich spectrum.
@@
In particular, $\mathfrak{a}_{\bullet}(v)$ is saturated for any $v\in\mathrm{Val}^{+}_{R,\mathfrak{m}}$. By the above proposition, saturated filtrations (or norms) can be considered as the local analogue of maximal norms in the sense of [BJ25, Definition 6.16].
@@
(2.4) $\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}}).$
@@
(2.5) $\mathfrak{a}_{\bullet}=\cap_{v\in\mathrm{DivVal}_{R,\mathfrak{m}}}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a}_{\bullet})})$
@@
applying the formula (2.5) to $\widetilde{\mathfrak{a}}_{\bullet}$ and $\widetilde{\mathfrak{b}}_{\bullet}$ respectively, we know that $\mathfrak{a}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}\subset\mathfrak{c}_{\bullet}$ and similarly $\mathfrak{b}_{\bullet}\subset\mathfrak{c}_{\bullet}$. Hence $v(\mathfrak{c}_{\bullet})\leq\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$ for any $v\in\mathrm{Val}_{R,\mathfrak{m}}$. By (2.4), for any $v\in\mathrm{DivVal}_{R,\mathfrak{m}}$ we have $v(\mathfrak{c}_{\bullet})\geq v(\mathfrak{a}_{\bullet}(v/\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}))=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}$. Thus
(2.6) $v(\mathfrak{c}_{\bullet})=\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\}.$
[…]
@@
$=$ $\max\{\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{b}_{\bullet})\},\min\{v(\mathfrak{a}_{\bullet}),v(\mathfrak{c}_{\bullet})\}\}$
@@
where the second inequality follows from Proposition 2.23, and the last inequality is easy to check. By Proposition 2.23 again, we get $f\in(\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet}))_{\lambda}$. This shows that $\mathfrak{a}_{\bullet}\vee_{s}(\mathfrak{b}_{\bullet}\cap\mathfrak{c}_{\bullet})\supset(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{b}_{\bullet})\cap(\mathfrak{a}_{\bullet}\vee_{s}\mathfrak{c}_{\bullet})$ and the equality must hold. ∎
@@
[…]
@@
[…]
@@
(2.7) $\widetilde{\mathfrak{a}^{\bullet}}=\cap_{v\in RV(\mathfrak{a})}\mathfrak{a}_{\bullet}(\frac{v}{v(\mathfrak{a})}),$
@@ 2.3. Homogeneous filtrations
(2) Choose a set $\{f_{1},\ldots,f_{r}\}$ of generators of $\mathfrak{m}$ and $D\in\mathbb{Z}_{>0}$ such that $\mathfrak{m}^{D}\subset\mathfrak{a}_{1}$. Then $f_{i}^{D}\in\mathfrak{a}_{1}$, and hence $f_{i}\in\mathfrak{a}_{1/D}$ by assumption. So $\mathfrak{m}\subset\mathfrak{a}_{1/D}$ and hence $\mathfrak{a}_{\lambda}=\mathfrak{m}$ for $0<\lambda<1/D$. The last assertion is straightforward by Definition-Lemma 2.8. ∎
@@
(c.f. [HS06, Definition 6.9.3] and [BFJ14, Section 4]) Given a norm $\chi$ on $R$, its homogenization $\widehat{\chi}:R\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ is defined by $\widehat{\chi}(f):=\lim_{k\to\infty}\frac{\chi(f^{k})}{k}$ for any $f\in R$.
[…]
The limit exists by Fekete’s lemma, see, for example, [CP24, Theorem 3.4]. Hence it is not hard to see that $\widehat{\mathfrak{a}}_{\bullet}$ is the smallest homogeneous filtration that contains $\mathfrak{a}_{\bullet}$. Moreover, the homogenization has the following property.
@@
[CP24] Let $\mathfrak{a}_{\bullet}$ be an $\mathfrak{m}$-filtration. Then
  1. (1)
    $\widehat{\mathfrak{a}}_{\bullet}$ is the unique largest $\mathfrak{m}$-filtration $\mathfrak{b}_{\bullet}$ such that $\widehat{\mathrm{ord}_{\mathfrak{a}_{\bullet}}}=\widehat{\mathrm{ord}_{\mathfrak{b}_{\bullet}}}$, and
  2. (2)
    $\widehat{\mathfrak{a}}_{\bullet}\subset\widetilde{\mathfrak{a}}_{\bullet}$.
Before proceeding, we briefly recall the definition of the Berkovich analytification. Fix $\chi\in\mathcal{N}$. By linear boundedness, we know that the $\mathfrak{m}$-adic completion $\hat{R}$ is complete with respect to $\chi$. Let $\mathcal{M}=\mathcal{M}(R,\chi)$ be the Berkovich spectrum of the Banach ring $(\hat{R},\chi)$, that is, the set of multiplicative $\mathfrak{m}$-seminorms $w$ on $\hat{R}$ such that $w\geq\chi$. Note that the homogenization corresponds to the spectral radius construction of [Ber90], since we adopted the multiplicative notation. We thank the referee for pointing out a mistake in the statement of the previous version.
Next, we give a characterization for homogeneous norms similar to Proposition 2.20, which is an easier local analogue of [Fin25b, Theorem 3.17] and [BJ22, Theorem 2.16].
@@
$\chi=\inf_{w\in\Sigma}w,$
where $\Sigma\subset\mathcal{M}$ is a non-empty subset.
@@
If $\chi(f)=\inf_{w\in I}w(f)$, then $\chi$ is homogeneous, since the same is true for each $w$.
The reverse implication is essentially a restatement of the Berkovich Maximum Modulus principle, see [Ber90, Theorem 1.3.1]. ∎
@@
For adic filtrations, it is not hard to see that homogenization and saturation coincide.
Lemma 2.33.
Let $\mathfrak{a}\in\mathrm{Fil}_{R,\mathfrak{m}}$. Then $\widetilde{\mathfrak{a}^{\bullet}}=\widehat{\mathfrak{a}^{\bullet}}$.
Proof.
By Example 2.25, $(\widetilde{\mathfrak{a}^{\bullet}})_{m}=\overline{\mathfrak{a}^{m}}$ is the integral closure of $\mathfrak{a}^{m}$ for any $m\in\mathbb{Z}_{>0}$. The same calculation implies $(\widehat{\mathfrak{a}^{\bullet}})_{m}=\overline{\mathfrak{a}^{m}}$ by [HS06, Proposition 6.10]. So the lemma follows from a simple calculation since both filtrations are homogeneous by Lemma 2.29. ∎